# VaLID: Verification as Late Integration of Detections for LiDAR-Camera Fusion

### v1

@@ VI Limitations
Our late fusion method is targeted at reducing false positives, but does not seek to increase true positives. We thus require that the LiDAR (or primary) detector successfully finds most true positives. Similarly, since our approach relies on validating LiDAR detections with camera inputs, it may not perform well with camera models that produce few detections, and is therefore more applicable to methods which can be tuned for a greater number.
@@ VII Conclusion
In this paper, we propose a model-independent late fusion method that combines LiDAR and camera detections to enhance object detection performance for autonomous vehicles. Our approach focuses on reducing false positives by cross-validating LiDAR detections with camera inputs. Experiments on the KITTI dataset demonstrate that our method is exceptional at reducing false positives, and performs competitively with state-of-the-art methods on average precision. Importantly, the method works equally well with several camera detectors, including non-specialized and open-vocabulary models.

### v2

@@ VI Limitations
Our late fusion method is targeted at reducing false positives, but does not seek to increase true positives. We thus require that the LiDAR (or primary) detector successfully finds most true positives. Similarly, since our approach relies on validating LiDAR detections with camera inputs, it may not perform well with camera models that produce few detections and is, therefore, more applicable to methods that can be tuned for a greater number.
@@ VII Conclusion
In this paper, we propose a model-adaptable late fusion method that combines LiDAR and camera detections to enhance object detection performance for autonomous vehicles. Our approach focuses on reducing false positives by cross-validating LiDAR detections with camera inputs. Experiments on the KITTI dataset demonstrate that our method is exceptional at its primary goal of reducing false positives, and provides modest but consistent improvements in average precision. The method performs competitively with state-of-the-art comparison methods.
Importantly, the method works equally well with relatively low precision general purpose camera detectors which were not specifically trained on this dataset or for these specific detection classes. This is a surprising result, which we hope sparks future research on using general-purpose 2D detectors to improve the special purpose 3D vehicle detectors which currently dominate the state of the art.
Acknowledgments
Partial support to some authors for this work was provided by WISEautomotive through the ATC+ Program award from MOTIE Korea 20014264.